OpenStudy (abmon98):

The equation of a curve is xy=12 and the equation of a line L is 2x+y=k where k is a constant. I) in the case k=11 find the coordinates of the points of intersection of L and the curve. Ii) find the set of values of k for which L does not intersect the curve. Iii) in the case where k=10 one of the points of intersection is p(2,6). Find the angle in degrees between L and the tangent to the curve at P

OpenStudy (anonymous):

y = 12 / x 2x + y = 11 2x + ( 12 / x ) = 11 2x^2 - 11x + 12 = 0 (2x - 3 ) ( x - 4 ) = 0 x = 3 /2, y = 8 or x = 4, y = 3

OpenStudy (abmon98):

I solves I but iam stuck on both ii and iii and thank you

OpenStudy (abmon98):

Solved i*

OpenStudy (abmon98):

Should I differentiate

OpenStudy (abmon98):

And k is not equal to 11

OpenStudy (abmon98):

When I looked back to my textbook answer appears to be 9.8 and -9.8

OpenStudy (loser66):

ii) y = 12/x 2x+y =k replace y to second one you have 2x^2 +12-kx =0 discriminant <0 --> no solution so (-k)^2 -4*2*12 <0 k^2 <96 -9.8< k <9.8

OpenStudy (loser66):

a little bit faster, :)

OpenStudy (abmon98):

Thank you so so much but why is m=3

OpenStudy (abmon98):

Oh now I understand thank you man

OpenStudy (loser66):

https://www.desmos.com/calculator